# latexpage numbers missing

### v1

@@ 2.4 Discretization
A popular SDE used for score-based diffusion models is the VPSDE, which, in the conditional setting, sets $\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)=\frac{-\beta(t)}{2}\tilde{\bm{x}}_{t}$ and $g(t)=\sqrt{\beta(t)}$ in (5b), where $\beta(t):[0,T]\to\mathbb{R}^{+}$ is selected in such a manner that the process has bounded variance as $T\to\infty$ (Song et al., 2021). To evolve $\tilde{\bm{x}}_{t}$ with the conditional forward and reverse SDEs, we discretize time. The discretized versions of (5b) and (6b) are given by
@@
where $\beta_{t}=\beta(t)$ and $\tilde{\bm{\epsilon}}_{t}$ is defined on $\tilde{\mathcal{S}}$. See Supplementary Materials S1.2 and Appendix B in Song et al. (2021) for derivation details. The corresponding transition kernels in the discretized conditional forward SDE have the form
@@
where $\underline{\sigma}_{t}=1-\underline{\alpha}_{t}$, $\underline{\alpha}_{t}=\prod_{s=1}^{t-1}\alpha_{s}$ and $\alpha_{s}=1-\beta_{s}$. To sample from the target distribution $p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0})$ using (9), we replace the true score function $\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))$ with the approximation $\tilde{\bm{s}}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},t)$. Note that in (7), $t$ now needs to be equipped with a probability mass function on $\{1,\dots,T\}$ instead of a continuous one. The VPSDE framework for score-based diffusion in (8) and (9) is sometimes also referred to as a denoising diffusion probabilistic model (DDPM); Chapter 4 in Chan (2025) clarifies the connection between these two models.
@@ 2.5 Architecture for Conditional Score Neural Network
A key consideration in our methodology is the architecture of the neural network surrogate for the conditional score function $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ which must be flexible enough to accurately approximate highly non-linear score functions. Here, we restrict $\mathcal{S}$ to points on a regular grid, which enables the use of convolutional neural network (CNN) architectures (LeCun et al., 1998). In what follows, we regard each realization of the spatial process on $\mathcal{S}$ as an image, and $\tilde{\mathcal{S}}$ corresponds to pixels within the image that are treated as unobserved. To cater to score functions of variable dimension, we train a neural network whose input and output at time $t\in[0,T]$ has dimension $\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and which outputs 0 for all observed locations. That is, our neural network,
$\bm{s}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t)=(\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t),\bm{\mathring{0}}_{\mathring{\mathcal{S}}})\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$, (11)
yields the appropriate subvector $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t)$ for use in (7) and (9) as part of its output. For brevity, we assume (11) places the values at the correct locations to avoid cumbersome notation.
For $\bm{s}_{\bm{\phi}}(\cdot,\cdot,\cdot)$, we use a specific type of CNN, referred to as U-Net, which consists of two main parts: a contracting path (or encoder) which reduces the image dimension and yields a large number of feature maps through multiple layers, and an expanding path (or decoder), that uses several layers of transposed convolutions, skip connections, and upsampling stages to restore the original spatial resolution; see Ronneberger et al. (2015) for more details. There are many variants of U-Nets. In our implementation, we employ the noise conditional score network++ model (NCSN++) of Song et al. (2021) with minor modifications to reduce the U-Net size and thus, decrease training and evaluation time. Specifically, we reduced the number of blocks in both the contracting and expanding paths of the U-Net from $4$ to $2$ and the number of Gaussian Fourier projections determining the sinusoidal time embeddings from $128$ to $64$.
Recall from (7) that the conditional score depends on the observed locations, and therefore $\mathring{\mathcal{S}}$ needs to be an input to the score matching neural network. Since CNNs can only handle inputs of fixed dimension, we instead input a mask $\bm{M}(\mathring{\mathcal{S}})\in\{0,1\}^{\lvert S\rvert}$ where
@@ 2.6 Validation
Validating whether the neural conditional simulations come from the true predictive distributions is difficult due to the high-dimensional setting, the assumed intractability of the predictive distributions, as well as the variable conditioning set. To address these issues, we propose a sample-based validation method derived from Linhart et al. (2023), which compares the NCS-approximated unconditional distribution (described below) to the true unconditional distribution. This approach has multiple benefits: First, this method is tractable because simulation via NCS and from the true unconditional distribution is computationally efficient. Second, the NCS-approximated unconditional distribution does not depend on a single conditioning set, and thus, we can sidestep the issue of verifying the predictive distribution for each and every possible conditioning set. Finally, this validation technique enables the comparison of multivariate metrics, such as the spatial minimum and maximum of each unconditional simulation, which will only follow the correct distribution if the corresponding high-dimensional characteristics of the true unconditional distributions are accurately captured. However, as with any validation technique, passing these checks is a necessary but not sufficient condition for correct conditional simulation.
In addition to these multivariate metrics, we also use low-dimensional metrics. For the low-dimensional metrics, we created an evaluation data set involving replicates generated via NCS and either the true predictive distribution or an MCMC approximation of the true predictive distribution. Since these low-dimensional metrics are straightforward summaries, here we only discuss how to generate the NCS-approximated unconditional distribution in depth, and defer the technical details for the low-dimensional validation metrics to Supplementary Materials S2.3.
@@
Consider the setting in which parameters are fixed and the conditioning sets are drawn from a fixed distribution $p(\mathring{\mathcal{S}})$. We simulate $m$ replicates $\{\bm{x}_{0,i}\}_{i\in[m]}$ from the true unconditional distribution and $m$ corresponding conditioning sets $\{\mathring{\mathcal{S}}_{i}\}_{i\in[m]}$ and masks $\{\bm{M}(\mathring{\mathcal{S}}_{i})\}_{i\in[m]}$ according to $p(\mathring{\mathcal{S}})$. For each of the $m$ replicates $\bm{x}_{0,i}$ and masks $\bm{M}(\mathring{\mathcal{S}}_{0,i})$, we obtain the corresponding partially observed spatial fields $\bm{\mathring{x}}_{0,i}$, generate a single NCS simulation $\tilde{\bm{x}}_{0,i}$ conditional on $\bm{\mathring{x}}_{0,i}$, and concatenate the NCS simulation with the partially observed field to produce an “unconditional” simulation $\bm{\mathring{x}}_{0,i},\tilde{\bm{x}}_{0,i}]$. If NCS accurately samples from $p(\tilde{\bm{x}}_{0,i}\mid\bm{\mathring{x}}_{0,i})$, then ${\bm{\mathring{x}}}_{0,i},\tilde{\bm{x}}_{0,i}]$ is a sample from $p(\bm{x}_{0})$, and can therefore be validated against other samples drawn from $p(\bm{x}_{0})$. See Supplementary Materials S2.3 for pseudocode of this procedure.
@@ 3 Simulation Experiments
To demonstrate NCS, we use it to conditionally simulate from two spatial processes: a Gaussian process and a Brown–Resnick process. The first Gaussian case study verifies that our method is able to generate samples from true (known) predictive distributions; the second Brown–Resnick case study serves to demonstrate the utility of our method in settings where fast and efficient unconditional simulation is possible, but from which the predictive distributions are difficult or intractable to sample. For brevity, here we only present the results for the Brown–Resnick case study; see Supplementary Materials S3 for the Gaussian case study.
@@ 3.1 Experimental Setup
We let the spatial locations in $\mathcal{S}$ be a 32 $\times$ 32 gridding on $\mathcal{D}=[-10,10]\times[-10,10]$. The conditioning sets $\mathring{\mathcal{S}}$ and corresponding masks $\bm{M}(\mathring{\mathcal{S}})$ were constructed either by sampling at random a fixed number of locations from the $32^{2}$ grid points (Section 3.2), or through $32^{2}$ independent Bernoulli draws with probability of success (corresponding to whether a pixel is observed or not) equal to $\rho$ (Section 3.3). For the diffusion processes in (8) and (9), we set $T=1000$, $\beta_{0}=0.0001$, $\beta_{T}=0.02$ and we let $\beta_{t}=\beta(t)$ increase linearly with $t$ (specifically, we let $\beta(t)=0.0001+0.0199t/T$). We used a weighting function $\lambda(t)=1-\underline{\alpha}_{t}$ for $\underline{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ in the conditional loss function (7); see Song et al. (2021) and Supplementary Materials S2.2 for details. Training and validation data were generated “on the fly” (Chan et al., 2018).
@@
The Brown–Resnick process that we use in this paper is a two-parameter max-stable process with $\bm{\theta}\equiv(\lambda,\nu)$, with range parameter $\lambda\in\mathbb{R}^{+}$ and smoothness parameter $\nu\in(0,2]$. It admits a likelihood function that is extremely costly to evaluate, and predictive distributions that are computationally challenging to sample from (Padoan et al., 2010; Davison et al., 2012; Dombry et al., 2013; Huser et al., 2019). For the purpose of assessing NCS, we consider the Gibbs sampler (Dombry et al., 2013) implemented in the SpatialExtremes R package (Ribatet, 2020) as an alternative conditional simulation method. We found that full conditional simulation (FCS) via the SpatialExtremes package is only computationally feasible when the conditioning set contains seven or fewer observations. For conditioning sets with more than seven observed locations, we resort to local conditional simulation (LCS) by only using the the seven observed locations closest to a prediction location for conditional simulation. In contrast to FCS and NCS, LCS targets a low-dimensional approximation of a predictive distribution and will only be used here for univariate and bivariate approximations of the predictive distributions. The Gibbs sampler is computationally intensive, and requires a computing wall time that depends on several factors, such as the number of observed locations and model parameters. In contrast, as we show below in Section 3.2, NCS allows for consistent computational efficiency, regardless of the number of observed locations or process-model parameters. We implemented NCS and LCS/FCS on an Nvidia Tesla T4 GPU and an AMD EPYC 7V12 processor (with 64 cores, no multithreading) respectively.
For the Brown–Resnick process, we log-transformed the simulated spatial fields from the unit Fréchet scale to the standard Gumbel scale before training the U-Nets for estimating the score function. The learning rate, batch size, and number of epochs varied by U-Net and case study. See Supplementary Materials S2.2 for details.
@@ 3.2 Results: Small Number of Observed Locations
In this section, we compare FCS and NCS for the case when the number of observed locations $|\mathring{\mathcal{S}}|\in\{1,\dots,7\}$ (where FCS is computationally feasible) and where the range parameter $\lambda\in\{1,\dots,5\}$ (weak to strong spatial dependence) varies. We trained and implemented separate U-Nets for each observation/parameter combination. See Supplementary Materials S2.2 for details. To assess NCS, we examine the extremal correlation function $\chi(h)=2-\zeta(h)$, where the spatial extremal coefficient function $\zeta(h):\mathbb{R}^{+}\to[1,2]$ quantifies the extremal dependence between pairs of locations a fixed distance $h$ apart (de Haan, 1985; Smith, 1990). We compute the empirical extremal correlation functions for the true unconditional distributions, the FCS-approximated unconditional distributions, and the NCS-approximated unconditional distributions using the approach described in Section 2.6 (Linhart et al., 2023). In addition to the extremal correlation functions, we also empirically estimate the unconditional densities of the minimum, maximum, and the absolute sum of the process over $\mathcal{S}$. See Supplementary Materials S4.3 for details.
@@
To highlight the computational inefficiency and ultimately infeasibility of FCS when compared to NCS, we record the average time (in seconds) to generate one conditional simulation for varying number of observations, and divide it by the time needed to conditionally simulate with $|\mathring{\mathcal{S}}|=1$. As shown through these time ratios in Figure 2, the average time to generate an FCS simulation when seven locations are observed is five times greater than the time needed when one location is observed. Since the Gibbs sampler underlying FCS involves evaluating multivariate Gaussian cumulative distribution functions, we expect FCS to scale exponentially with the number of observed locations $|\mathring{\mathcal{S}}|$ (Hinrichs et al., 2014; Balakrishnan, 2014; Ribatet, 2020) and use this to show the likely scaling of FCS beyond seven observations—the maximum number of locations for which our implementation of the Gibbs sampler in the SpatialExtremes R package remained stable. We conclude that FCS is difficult to implement using available software, and its computational complexity quickly increases with the number of observed locations.
On the other hand, the time needed to generate an NCS simulation does not change with the number of observed locations. The average time to generate a single FCS and NCS simulation from seven observed locations was $28$ and $27$ seconds, respectively, with our setup. We currently evaluate the U-Net at every time step in the discretized conditional reverse SDE, but there are methods designed to reduce the number of steps needed to simulate from diffusion models (Bond-Taylor et al., 2022). Once implemented, these may further reduce the time needed for a single NCS simulation.

### v2

@@ 2.4 Discretization
A popular SDE used for score-based diffusion models is the VPSDE, which, in the conditional setting, sets $\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)=\frac{-\beta(t)}{2}\tilde{\bm{x}}_{t}$ and $g(t)=\sqrt{\beta(t)}$ in (5b), where $\beta(t):[0,T]\to\mathbb{R}^{+}$ is selected in such a manner that the process has bounded variance as $T\to\infty$ (Song2021). To evolve $\tilde{\bm{x}}_{t}$ with the conditional forward and reverse SDEs, we discretize time. The discretized versions of (5b) and (6b) are given by
@@
where $\beta_{t}=\beta(t)$ and $\tilde{\bm{\epsilon}}_{t}$ is defined on $\tilde{\mathcal{S}}$. See Supplementary Materials S1.2 and Appendix B in Song2021 for derivation details. The corresponding transition kernels in the discretized conditional forward SDE have the form
@@
where $\underline{\sigma}_{t}=1-\underline{\alpha}_{t}$, $\underline{\alpha}_{t}=\prod_{s=1}^{t-1}\alpha_{s}$ and $\alpha_{s}=1-\beta_{s}$. To sample from the target distribution $p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0})$ using (9), we replace the true score function $\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))$ with the approximation $\tilde{\bm{s}}_{\bm{\phi}}(\bm{x}_{t},\mathring{\mathcal{S}},t)$. Note that in (7a), $t$ now needs to be equipped with a probability mass function on $\{1,\dots,T\}$ instead of a continuous one. The VPSDE framework for score-based diffusion in (8) and (9) is sometimes also referred to as a denoising diffusion probabilistic model (DDPM); Chapter 4 in Chan2025 clarifies the connection between these two models.
@@ 2.5 Architecture for Conditional Score Neural Network
A key consideration in our methodology is the architecture of the neural network surrogate for the conditional score function $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ which must be flexible enough to accurately approximate highly non-linear score functions. Here, we restrict $\mathcal{S}$ to points on a regular grid, which enables the use of convolutional neural network (CNN) architectures (Lecun1998). In what follows, we regard each realization of the spatial process on $\mathcal{S}$ as an image, and $\tilde{\mathcal{S}}$ corresponds to pixels within the image that are treated as unobserved. To cater to score functions of variable dimension, we train a neural network whose input and output at time $t\in[0,T]$ has dimension $\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and which outputs 0 for all observed locations. That is, our neural network,
$\bm{s}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t)=(\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t)^{\prime},\bm{\mathring{0}}_{\mathring{\mathcal{S}}}^{\prime})^{\prime}\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$, (11)
yields the appropriate subvector $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\mathring{\mathcal{S}},t)$ for use in (7a) and (9) as part of its output. For brevity, we assume (11) places the values at the correct locations to avoid cumbersome notation.
For $\bm{s}_{\bm{\phi}}(\cdot,\cdot,\cdot)$, we use a specific type of CNN, referred to as U-Net, which consists of two main parts: a contracting path (or encoder) which reduces the image dimension and yields a large number of feature maps through multiple layers, and an expanding path (or decoder), that uses several layers of transposed convolutions, skip connections, and upsampling stages to restore the original spatial resolution; see Ronneberger2015 for more details. There are many variants of U-Nets. In our implementation, we employ the noise conditional score network++ model (NCSN++) of Song2021 with minor modifications to reduce the U-Net size and thus, decrease training and evaluation time. Specifically, we reduced the number of blocks in both the contracting and expanding paths of the U-Net from $4$ to $2$ and the number of Gaussian Fourier projections determining the sinusoidal time embeddings from $128$ to $64$.
Recall from (7a) that the conditional score depends on the observed locations, and therefore $\mathring{\mathcal{S}}$ needs to be an input to the score matching neural network. Since CNNs can only handle inputs of fixed dimension, we instead input a mask $\bm{M}(\mathring{\mathcal{S}})\in\{0,1\}^{\lvert S\rvert}$ where
@@ 2.6 Validation
Validating whether the neural conditional simulations come from the true predictive distributions is difficult due to the high-dimensional setting, the assumed intractability of the predictive distributions, as well as the variable conditioning set. To address these issues, we propose a sample-based validation method derived from Linhart2023, which compares the NCS-approximated unconditional distribution (described below) to the true unconditional distribution. This approach has multiple benefits: First, this method is tractable because simulation via NCS and from the true unconditional distribution is computationally efficient. Second, the NCS-approximated unconditional distribution does not depend on a single conditioning set, and thus, we can sidestep the issue of verifying the predictive distribution for each and every possible conditioning set. Finally, this validation technique enables the comparison of multivariate metrics, such as the spatial minimum and maximum of each unconditional simulation, which will only follow the correct distribution if the corresponding high-dimensional characteristics of the true unconditional distributions are accurately captured. However, as with any validation technique, passing these checks is a necessary but not sufficient condition for correct conditional simulation.
In addition to these multivariate metrics, we also use low-dimensional metrics. For the low-dimensional metrics, we created a validation data set involving replicates generated via NCS and either the true predictive distribution or an MCMC approximation of the true predictive distribution. Since these low-dimensional metrics are straightforward summaries, here we only discuss how to generate the NCS-approximated unconditional distribution in depth, and defer the technical details for the low-dimensional validation metrics to Supplementary Materials S2.3.1.
@@
Consider the setting in which parameters are fixed and the conditioning sets are drawn from a fixed distribution $p(\mathring{\mathcal{S}})$. We simulate $m$ replicates $\{\bm{x}_{0,i}\}_{i\in[m]}$ from the true unconditional distribution and $m$ corresponding conditioning sets $\{\mathring{\mathcal{S}}_{i}\}_{i\in[m]}$ and masks $\{\bm{M}(\mathring{\mathcal{S}}_{i})\}_{i\in[m]}$ according to $p(\mathring{\mathcal{S}})$. For each of the $m$ replicates $\bm{x}_{0,i}$ and masks $\bm{M}(\mathring{\mathcal{S}}_{0,i})$, we obtain the corresponding partially observed spatial fields $\bm{\mathring{x}}_{0,i}$, generate a single NCS simulation $\tilde{\bm{x}}_{0,i}$ conditional on $\bm{\mathring{x}}_{0,i}$, and concatenate the NCS simulation with the partially observed field to produce an “unconditional” simulation $(\bm{\mathring{x}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}$. If NCS accurately samples from $p(\tilde{\bm{x}}_{0,i}\mid\bm{\mathring{x}}_{0,i})$, then $({\bm{\mathring{x}}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}$ is a sample from $p(\bm{x}_{0})$, and can therefore be validated against other samples drawn from $p(\bm{x}_{0})$. See Algorithms S3 and S4 in the Supplementary Materials for pseudocode of this procedure.
@@ 3 Simulation Experiments
To demonstrate the accuracy of NCS, we use it to conditionally simulate from two spatial processes: a Gaussian process and a Brown–Resnick process. The first Gaussian case study verifies that our method is able to generate samples from true (known) predictive distributions; the second Brown–Resnick case study serves to demonstrate the utility of our method in settings where fast and efficient unconditional simulation is possible, but from which the predictive distributions are difficult or intractable to sample. For brevity, here we only present the results for the Brown–Resnick case study; see Supplementary Materials S3 for the Gaussian case study.
@@ 3.1 Experimental Setup
We let the spatial locations in $\mathcal{S}$ be a 32 $\times$ 32 gridding on $\mathcal{D}=[-10,10]\times[-10,10]$. The conditioning sets $\mathring{\mathcal{S}}$ and corresponding masks $\bm{M}(\mathring{\mathcal{S}})$ were constructed either by sampling at random a fixed number of locations from the $32^{2}$ grid points (Section 3.2), or through $32^{2}$ independent Bernoulli draws with probability of success (corresponding to whether a pixel is observed or not) equal to $\rho$ (Section 3.3). For the diffusion processes in (8) and (9), we set $T=1000$, $\beta_{0}=0.0001$, $\beta_{T}=0.02$ and we let $\beta_{t}=\beta(t)$ increase linearly with $t$ (specifically, we let $\beta(t)=0.0001+0.0199t/T$). We used a weighting function $\lambda(t)=1-\underline{\alpha}_{t}$ for $\underline{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ in the conditional loss function (7a); see Song2021 and Supplementary Materials S2.2 and S4.2 for details. Training and validation data were generated “on the fly” (Chan2018).
The Brown–Resnick process that we use in this paper is a two-parameter max-stable process with $\bm{\theta}\equiv(\lambda,\nu)^{\prime}$, with range parameter $\lambda\in\mathbb{R}^{+}$ and smoothness parameter $\nu\in(0,2]$. It admits a likelihood function that is extremely costly to evaluate, and predictive distributions that are computationally challenging to sample from (Padoan2010; Davison; Dombry2012; Huser2019). See Supplementary Materials S2.1 for greater detail on the Brown–Resnick process.
@@
For the purpose of assessing NCS, we consider the Gibbs sampler (Dombry2012) implemented in the SpatialExtremes R package (SpatialExtremes) as an alternative conditional simulation method. We found that full conditional simulation (FCS) via the SpatialExtremes package is only computationally feasible when the conditioning set contains seven or fewer observations. For conditioning sets with more than seven observed locations, we resort to local conditional simulation (LCS) by only using the the seven observed locations closest to a prediction location for conditional simulation. In contrast to FCS and NCS, LCS targets a low-dimensional approximation of a predictive distribution and will only be used here for univariate and bivariate approximations of the predictive distributions. The Gibbs sampler is computationally intensive, and requires a computing wall time that depends on several factors, such as the number of observed locations and model parameters. See Supplementary Materials S4.1 for more details.
In contrast, as we show below in Section 3.2, NCS allows for consistent computational efficiency, regardless of the number of observed locations or process-model parameters. We implemented NCS and LCS/FCS on an Nvidia Tesla T4 GPU and an AMD EPYC 7V12 processor (with 64 cores, no multithreading) respectively. For the Brown–Resnick process, we log-transformed the simulated spatial fields from the unit Fréchet scale to the standard Gumbel scale before training the U-Nets for estimating the score function.
@@ 3.2 Results: Small Number of Observed Locations
In this section, we compare FCS and NCS for the case when the number of observed locations $|\mathring{\mathcal{S}}|\in\{1,\dots,7\}$ (where FCS is computationally feasible) and where the range parameter $\lambda\in\{1,\dots,5\}$ (weak to strong spatial dependence) varies. For each range value, we trained and implemented separate U-Nets that are each amortized for conditioning sets of $1$ to $7$ locations. See Supplementary Materials S2.2 for details. To assess NCS, we examine the extremal correlation function $\chi(h)=2-\zeta(h)$, where the spatial extremal coefficient function $\zeta(h):\mathbb{R}^{+}\to[1,2]$ quantifies the extremal dependence between pairs of locations a fixed distance $h$ apart (deHaan1985; Smith1990). We compute the empirical extremal correlation functions for the true unconditional distributions, the FCS-approximated unconditional distributions, and the NCS-approximated unconditional distributions using the approach described in Section 2.6 (Linhart2023). In addition to the extremal correlation functions, we also empirically estimate the unconditional densities of the minimum, maximum, and the absolute sum of the process over $\mathcal{S}$. See Supplementary Materials S4 for training and validation details as well as additional results.
@@
To highlight the computational inefficiency and ultimately infeasibility of FCS when compared to NCS, we record the average time (in seconds) to generate one conditional simulation for varying number of observations, and divide it by the time needed to conditionally simulate with $|\mathring{\mathcal{S}}|=1$. As shown through these time ratios in Figure 2, the average time to generate an FCS simulation when seven locations are observed is five times greater than the time needed when one location is observed. Since the Gibbs sampler underlying FCS involves evaluating multivariate Gaussian cumulative distribution functions, we expect FCS to scale exponentially with the number of observed locations $|\mathring{\mathcal{S}}|$ (Hinrich2014; Balakrishnan2014; SpatialExtremes) and use this to show the likely scaling of FCS beyond seven observations—the maximum number of locations for which our implementation of the Gibbs sampler in the SpatialExtremes R package remained stable. We conclude that FCS is difficult to implement using available software, and its computational complexity quickly increases with the number of observed locations.
On the other hand, the time needed to generate an NCS simulation does not change with the number of observed locations. The average time to generate a single FCS and NCS simulation from seven observed locations was $28$ and $27$ seconds, respectively, with our setup. We currently evaluate the U-Net at every time step in the discretized conditional reverse SDE, but there are methods designed to reduce the number of steps needed to simulate from diffusion models (Bond-Taylor2022). Once implemented, these may further reduce the time needed for a single NCS simulation.